OpenStudy (anonymous):

Solve the equation: u′′+4u+5u=0 With the initital values: u(0)=1 u′(0)=0

OpenStudy (angela210793):

u′′ wht's this? O.o

OpenStudy (anonymous):

d^2y/dx^2

OpenStudy (anonymous):

Aux equation is: \[\lambda^{2}+4 lambda + 5=0\]

OpenStudy (anonymous):

\[\lambda = - 2 \pm i\]

OpenStudy (anonymous):

Complex conjugate roots, keep going..

OpenStudy (anonymous):

Woops made a mistake in the initital equation should be: \[u''+4u'+5u=0\]

OpenStudy (anonymous):

I guessed that...:-

OpenStudy (anonymous):

So general solution is...?

OpenStudy (anonymous):

Then the general equation is: \[u = Ae ^{\lambda}+Be^{\lceil lamdba \rceil}\] no complex conjugate symbol so I used ceiling

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[e^{-2t}(Acos(t)+Bsin(t))\]

OpenStudy (anonymous):

thats general soln

OpenStudy (anonymous):

then its just a ton of product rules with the conditions to find the constants

OpenStudy (anonymous):

which can be written: \[u=Ae^{\alpha x}e^{i \beta x} + Be^{\alpha x}e^{-i \beta x}\]

OpenStudy (anonymous):

so finding it for u I just plug in zero

OpenStudy (anonymous):

Yes, elecengineer has just substituted in already...

OpenStudy (anonymous):

and make it equal to 0

OpenStudy (anonymous):

Think you have gone off the track. Go back to where elecengineer put up the general solution (in t instead of u).

OpenStudy (anonymous):

assuming u as function of t general soln \[u(t) = e^{-2t}(Acos(t) +Bsin(t) )\]

OpenStudy (anonymous):

u(0) = 1 gives 1= 1( A) ie A=1

OpenStudy (anonymous):

then differenitate du/dt = e^(-2t) ( -sin(t) +Bcos(t) ) +(-2e^(-2t) ) ( cos(t) +Bsin(t) )

OpenStudy (anonymous):

\[\frac{du}{dt} = e^{-2t} ( -Asin(t) +Bcos(t) ) -2e^{-2t}(Acos(t) +Bsin(t) ) \]

OpenStudy (anonymous):

sub A=1, du/dt =0 , t=0

OpenStudy (anonymous):

0 = 1(0+B) - 2(1)(1)

OpenStudy (anonymous):

B = 2

OpenStudy (anonymous):

B=2 soln is \[u(t) = e^{-2t} ( \cos(t) +2\sin(t))\]

OpenStudy (anonymous):

Sweet thanks

OpenStudy (anonymous):

Really helped with my revision